OpenStudy (anonymous):

Tony worksat a bike store. tony earns $300.00 every week plust $15.00 for every bike that he sells. write an inequality that can be used to determine the number of bike (b) tony must sell in one week if he wants to earn a minimum of $500.00 for that week. What is the minimum number of bikes tony must sell in one week to earn a weekly salary of $500.00?

OpenStudy (anonymous):

\[$ 300 + 15 \times n \ge $ 500\] \[n \ge \frac{$ 200}{15}\]

OpenStudy (anonymous):

thanks